OpenStudy (surryyy):

The area of a triangle is 75 square feet. If the height is 5 feet longer than the base, find the dimension of the height. Answer 10 15 20 25

OpenStudy (dumbcow):

area = 1/2(base*height)

OpenStudy (anonymous):

just take an answer minus it by 5 and multiply the number you get with the answer and apply it to the formula dumbcow put up there

OpenStudy (dumbcow):

---> 75 = 1/2( (height-5)*height) solve for height

OpenStudy (anonymous):

yea there

OpenStudy (dumbcow):

you have to use factoring a quadratic

OpenStudy (dumbcow):

Let x = height --> x(x-5) = 150 x^2 -5x -150 = 0 solve by factoring what 2 num add to -5 but mult to -150 ??

OpenStudy (surryyy):

would my final answer be 15?

OpenStudy (dumbcow):

that is correct :)

OpenStudy (dumbcow):

x^2 - 5x -150 = (x-15)(x+10) = 0 x = 15 or -10 height must be positive therefore x=15

OpenStudy (surryyy):

ok thnx!

OpenStudy (dumbcow):

yw